The table below shows how much output a firm is able to produce depending on how many resources it uses. The selling price for the output is $5 per unit, the cost of labor is $40 per unit, and the cost of capital is $15 per unit.

Labor

Output

Capital

MPL

Total Revenue

VMPL

0

0

3

1

24

3

2

44

3

3

60

3

4

72

3

5

80

3

6

84

3
Answer #1

Answer

the marginal product of n th worker =total output of n workers - the total output of n-1 workers
MPL(n)=TP(n)-TP(n-1)
MPL(1)=24-0=24

MPL(2)=44-24=20 and so on

=========

Total revenue =P*Q=output * price=5Q

TR(0)=0*5=0

TR(24)=24*5=120 and so on

===========

VMPL=MPL*Price

VMPL(1)=24*5=120

VMPL(2)=20*5=100 and so on

=============

Labor Output Capital MPL Total Revenue VMPL
0 0 3 0
1 24 3 24 120 120
2 44 3 20 220 100
3 60 3 16 300 80
4 72 3 12 360 60
5 80 3 8 400 40
6 84 3 4 420 20
6 84 3 0 420 0